Circular No. 6328

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

NOVALIKE VARIABLE IN SAGITTARIUS H. W. Duerbeck, European Southern Observatory, reports: “A 10-min

direct CCD image of Sakurai’s object (cf. IAUC 6322, 6323, 6325), taken on Feb. 29.4 UT with the 0.91-m Dutch telescope at La Silla through an Hα filter, shows that the star is surrounded by an almost-circular nebula of diameter 30′′; the strongest emission originates in the northwest quadrant. An exposure through a filter centered at 665 nm does not show any nebulosity, excluding the possibility that the shell is composed of illuminated dust. Thus Sakurai’s object is surrounded by a faint planetary nebula, a finding that is expected for an object undergoing a final helium flash (cf. IAUC 6325).”

D. Pollacco, Royal Greenwich Observatory, communicates: “Observations of this object by F. Verbunt, G. Geertsema, and M. van den Berg, using the 4.2-m William Herschel Telescope (+ ISIS spectrograph) on Feb. 29.25 UT, show it to be the central star of a limb-brightened planetary nebula (dimensions at least 45′′ north-south). The spectrum is typical of old, low-moderate excitation planetary nebulae with the [O III] 500.7nm/Hβ ratio ∼ 5 (very similar to the planetary nebula A58 surrounding the only other observed shell flash object, V605 Aql, in 1919). In view of the rareness of these events — this is the first one that can be studied with modern equipment — observers are strongly encouraged to monitor this object at both optical and infrared wavelengths. Spectroscopy over the coming months (years?) will be especially important as the material ejected during the shell flash is shocked/photoionized by the rapidly evolving star. Photometry may indicate massive dust ejection episodes, causing optical variations similar to those observed in R CrB stars. Infrared spectroscopy will first reveal the exposed Wolf-Rayet core.”

COMET C/1996 B2 (HYAKUTAKE) Total (visual) magnitude and coma-diameter estimates: Feb. 24.46 UT,

6.6, 18′ (C. S. Morris, Lockwood Valley, CA, 10×50 binoculars; 30′ tail in p.a. 278^{o}); 25.52, 6.9, 12′ (R. Keen, Mt. Thorodin, CO, 7×50 binoculars); 27.15, 6.7, ∼ 22′ (H. Mikuž, Ljubljana, Slovenia, 0.20-m reflector + CCD + V filter; faint, diffuse tail ∼ 1^{o} long in p.a. ∼ 290^{o}); 28.18, 6.4, – (J. D. Shanklin, Minera, Wales, 7×50 binoculars; 1^{o}.8 tail in p.a. 305^{o}); 29.37, 6.1, 15′ (G. W. Kronk, Troy, IL, 20×80 binoculars).

1996 February 29  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green